# OneVision: An End-to-End Generative Framework for Multi-view E-commerce Vision Search

### v1

@@ 1. Introduction
Recent years have witnessed substantial progress in vision search, propelled by advances across multiple fronts, including object detection (Cheng et al., 2024; Tian et al., 2025; Wang et al., 2025a), category prediction (Zhu et al., 2022; Noor et al., 2023, 2024), and visual representation learning (Lee et al., 2022a; Shao et al., 2023; Liang et al., 2025; Zhai et al., 2019; Siméoni et al., 2025). These developments have driven the construction of large-scale industrial systems, with major e-commerce platforms such as eBay (Yang et al., 2017) and Alibaba (Zhang et al., 2018; Chen et al., 2024a) deploying comprehensive vision search pipelines to support production-scale applications. Despite these advances, traditional pipelines in vision search still face structural limitations: (i) Difficulty in handling multi-view discrepancy. Query images and product images often differ significantly in viewpoint, making it hard for conventional retrieval pipelines to align their visual representations and recall the correct products. (ii) Stage-wise objective misalignment. The optimization goals across four stages collide, making it difficult to reach Pareto optimality in both user experience and conversion. (iii) Heavy multi-stage overhead. The multi-stage architecture introduces substantial computational cost and cross-module communication, increasing latency and complicating system maintenance.
Over the past two years, generative retrieval (GR) has emerged as a new paradigm for large-scale search, reframing retrieval as a sequence-to-sequence generation. Instead of relying on multi-stage matching, GR models directly generate discrete item identifiers from a query or user-behaviour sequence, thereby reducing pipeline complexity and alleviating the limitations of traditional retrieval architectures. OneRec (Deng et al., 2025) delivers a single-stage industrial recommender that unifies recall, pre-ranking, and ranking in one model. In e-commerce search, OneSearch (Chen et al., 2025) provides the first end-to-end generative retrieval framework with superior performance for high-quality recall and ranking. For image retrieval specifically, IRGen (Zhang et al., 2024) formulates image retrieval as seq2seq generation of nearest-neighbor IDs. GENIUS (Kim et al., 2025) adopts modality-decoupled semantic quantization and a query augmentation to enhance generalization. However, IRGen and GENIUS overlook multi-view discrepancies of the same object and ignore user-specific information, limiting their effectiveness in industrial vision search applications.
@@
3) Dynamic pruning for efficient inference. To address redundancy and inefficiency in image tokens, dynamic pruning techniques are applied to retain only the most informative visual tokens. In detail, visual tokens are compressed after the encoder using k-means clustering, and performance degradation is further mitigated through a curriculum learning and distillation framework. Inference efficiency is effectively improved with minimal impact on prediction accuracy.
@@
Extensive offline evaluations and ablation studies show that OneVision matches the performance of the online MCA, while dynamic pruning boosts inference efficiency by 21%. Large-scale online A/B tests on Kuaishou’s search platform further reveal notable improvements: CTR +2.15%, CVR +2.27%, OPM +3.60%, Clicks +1.80%, and Orders +3.12%. It demonstrates that OneVision effectively unifies retrieval and personalization, delivering measurable gains while simplifying the traditional serving pipeline.
@@ 2.2. Generative Retrieval System
Generative Retrieval application. Recent advances show that GR models can replace traditional recall–ranking stacks. OneRec (Deng et al., 2025) unifies recall and ranking via session-wise generation and DPO, achieving stable online gains with simpler serving. MMQ (Xu et al., 2025) further learns semantic IDs from multimodal content through a shared–specific tokenizer and behavior-aware fine-tuning to align semantics with user preferences. In e-commerce search, OneSearch (Chen et al., 2025) builds an end-to-end system where a keyword-enhanced hierarchical quantization encoder generates semantic IDs combining content, context, and collaborative signals; multi-view user sequences further improve personalization. These advances show GR can unify fragmented objectives, introduce personalization early, and reduce reliance on hand-crafted indices.
@@ 3.1. Vision-aligned Residual Quantization
A recent approach, RQ-OPQ (Chen et al., 2025), adopts RQ-KMeans for the shallow layers encoding. However, for vision search, this non-trainable clustering approach still struggles to ensure encoding consistency and uniqueness, especially for multi-view images of the same object. To address these issues, we propose VRQ, which integrates multi-view contrastive learning for trainable shallow codebooks and leverages statistical business features in residual OPQ encoding to enhance consistency and uniqueness. Additionally, we incorporate category information to suppress noisy samples and improve category-code consistency across the e-commerce image pool.
@@ 3.1.1. Collaborative Image Feature Alignment
In generative vision search systems, aligning similar product images or multi-view images enables the generation of high-quality visual representations, which are crucial for semantic encoding. Firstly, we construct a high-quality dataset of image pairs derived from real e-commerce user search logs, including both query–item and item–item pairs, covering multi-view images of the same product. Then, we apply cosine similarity for preliminary filtering, employ Qwen-VL (Bai et al., 2023) to further verify semantic consistency, and finally perform manual inspection and refinement to remove noisy data.
@@ 3.1.3. VRQ Hierarchical Quantization Encoding
Concretely, given a single product image and its associated category label, we first extract the visual feature $x$ using the image encoder $\mathcal{V}$ and the category feature $y$ using a pretrained BGE model (Xiao et al., 2024). We then learn latent representations $v:=\mathcal{E}_{v}(x)$ and $t:=\mathcal{E}_{t}(y)$, where $\mathcal{E}_{v}$ and $\mathcal{E}_{t}$ are modality-specific DNN encoders for the visual and textual inputs, respectively. To effectively combine multi-modal features into a unified representation, following (Baldrati et al., 2023), we compute the fused embedding as
@@
where $f_{i}^{(1)}$ and $f_{i}^{(2)}$ denote fused embeddings of the $i$-th positive product pair, $N$ is the number of positive pairs in a batch, and $\tau$ is a temperature parameter. To further enhance the integration of category information, we introduce a margin loss:
(4) $\mathcal{L}_{\mathrm{mar}}=\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{N}\left[\max\bigl(0,\ -\gamma+s_{ij}^{f2f}-s_{ii}^{v2f}\bigr)\right],$
@@ 3.2.1. Pretraining for Visual–Semantic Alignment
Considering that the basic architectures, such as BART and mT5, are pretrained on large-scale textual corpora, while the input of the vision search system is based on image representations, we first introduce a pretraining stage to establish initial semantic alignment among product images, titles, categories, and SIDs. To this end, we design four related tasks for pretraining: (i) Predicting the product SID from the corresponding image. (ii) Input both image and product title and output the corresponding SID. (iii) Predicting the product category from the image. (iv) Input both image and product title and output the corresponding category. The first task focuses on aligning visual representations with SIDs, while the remaining tasks introduce additional modalities to reinforce semantic relevance and facilitate alignment. During unified training, we insert a start token $t_{[BOS]}$ and an ending token $t_{[EOS]}$ at the first and last place, and use a separate token $t_{[SEP]}$ between different modalities. The training loss for the generative model is a cross-entropy loss for next-token prediction, formulated as:
@@ 3.2.3. Personalized Modeling and Post-training
To model user behaviour, we incorporate two complementary, per-user components: (i) user long-term behavioural sequence, represented as a single aggregated SID obtained by fusing the SIDs of multiple product the user has recently clicked or purchased on the Kuaishou e-commerce platform; and (ii) user short-term behavioural sequence, comprising the individual SIDs of the user’s five most recently clicked items during image-based search. In addition, we introduce explicit search-scene information to better distinguish intent across different contexts. The model is further trained with personalized SFT using the NTP loss to enhance its ability to capture nuanced user-specific purchasing preferences.
@@ 3.3. Dynamic Pruning
While the GR model offers strong semantic understanding, it also introduces significant computational overhead—particularly during beam search decoding. For a single query image, the model must perform $B$ rounds of decoding, where $B$ is the number of beam search candidates. This greatly impacts online inference speed. Moreover, the encoding of image tokens contains redundancy and inefficiency, both in the visual encoder and the generative model’s encoder. To mitigate this, we propose a dynamic pruning strategy that effectively retains only the most informative visual tokens. This significantly reduces inference latency with minimal impact on prediction accuracy.
@@ 3.4. End-to-End Inference
To produce a well-ranked candidate set, we use beam search. This method maintains multiple candidate SID sequences during generation, selecting them based on the accumulated log-probability scores. Starting with a $t_{[BOS]}$ token, the model predicts the top-$B$ most likely next tokens ($B$ being the beam size), and extends each sequence. After scoring the extended sequences, only the top-$B$ with the highest cumulative scores are retained. To prevent invalid SIDs, we apply constrained decoding using a Trie structure (De Cao et al., 2020; Fredkin, 1960). The Trie is built from the full set of valid SIDs before inference and restricts decoding to valid prefixes only. This ensures that generated sequences are valid and improves retrieval reliability.
@@ 4.1. Implementation Details
Datasets. We construct training datasets from large-scale industrial e-commerce search logs to support the multi-stage training of OneVision. The corpus consists of three parts: (1) query images and product images, each annotated with its corresponding SIDs, providing discrete representations for both search queries and products; (2) item-to-item (i2i) associations, representing semantically similar product pairs; and (3) 120 days’ data of user search logs from Kuaishou’s platform, covering item click and purchase records. Both the i2i and user-interaction data incorporate multi-view paired samples, enabling the model to build a more comprehensive and robust understanding of each product across diverse visual appearances. For evaluation, we built two test sets. The first is a real-world search interaction set, containing 8.5K click events and 8.5K purchase events sampled from 30 days of historical visual search traffic, designed to measure retrieval quality under actual user behaviours. The second is a semantic benchmark, comprising 67K query–item pairs (about 22K each from top, middle, and tail segments) collected via ANN recall. All datasets are additionally refined using ANN similarity scores and VLM verification to ensure reliable supervision and robust evaluation.
@@ 4.2. Real-World Offline Evaluation
Table 2 presents the ablation study of OneVision. Removing the ranking stage from onlineMCA leads to a noticeable decline in retrieval quality, confirming its importance. Substituting RQ–Kmeans with RQ–OPQ raises HR and MRR improves HR and MRR by 3.68% on average, showing that residual OPQ improves the stability of SIDs and enables a more fine-grained representation. The VRQ with multi-stage training reaches the highest overall scores while remaining comparable to the deployed online system. Adding statistical business characteristics (Convert) information brings an additional 2.98pt in HR and 3.22pt in MRR, which helps the model separate near-duplicate products more effectively. In contrast, user sequence modeling shows limited impact, likely because user preferences are sparse and relatively consistent, offering little additional signal for personalization. Category embeddings mainly mitigate category inconsistency and detection noise, lowering off-category retrievals and strengthening search robustness and user experience, while having a limited direct impact on HR or MRR.

### v3

@@ 1. Introduction
Recent years have witnessed substantial progress in vision search, propelled by advances across multiple fronts, including object detection (Cheng et al., 2024; Tian et al., 2025; Wang et al., 2025a), category prediction (Zhu et al., 2022; Noor et al., 2023, 2024), and visual representation learning (Lee et al., 2022a; Shao et al., 2023; Liang et al., 2025; Zhai et al., 2019; Siméoni et al., 2025). These developments have driven the construction of large-scale industrial systems, with major e-commerce platforms such as eBay (Yang et al., 2017) and Alibaba (Zhang et al., 2018; Chen et al., 2024a) deploying comprehensive vision search pipelines to support production-scale applications. Despite these advances, traditional pipelines in vision search still face structural limitations: (i) Difficulty in handling multi-view discrepancy. Query images and product images often differ significantly in viewpoint, making it hard for conventional retrieval pipelines to align their visual representations and recall the correct products. (ii) Stage-wise objective misalignment. The optimization goals across three stages collide, making it difficult to reach Pareto optimality in both user experience and conversion. (iii) Heavy multi-stage overhead. The multi-stage architecture introduces substantial computational cost and cross-module communication, increasing latency and complicating system maintenance.
Over the past two years, generative retrieval (GR) has emerged as a new paradigm for large-scale search, reframing retrieval as a sequence-to-sequence generation. Instead of relying on multi-stage matching, GR models directly generate discrete item identifiers from a query and user-behaviour sequence, thereby alleviating the limitations of traditional retrieval architectures. OneRec (Deng et al., 2025) delivers a single-stage industrial recommender that unifies recall, pre-ranking, and ranking in one model. In e-commerce search, OneSearch (Chen et al., 2025) provides the first end-to-end generative retrieval framework with superior performance for high-quality recall and ranking. For image retrieval specifically, IRGen (Zhang et al., 2024) formulates image retrieval as seq2seq generation of nearest-neighbor IDs. GENIUS (Kim et al., 2025) adopts modality-decoupled semantic quantization and a query augmentation to enhance generalization. However, IRGen and GENIUS overlook multi-view discrepancies of the same object and ignore user-specific information, limiting their effectiveness in industrial vision search applications.
@@
3) Dynamic pruning for efficient inference. To address redundancy and inefficiency in image tokens, dynamic pruning preserves only the most semantically informative visual tokens. Specifically, tokens are compressed before the decoder using K-means++ clustering, and potential performance degradation is mitigated via curriculum learning and distillation. This accelerates inference while maintaining accuracy and lowering compute cost.
@@
Extensive offline evaluations and ablation studies show that OneVision matches the performance of the online MCA, while dynamic pruning boosts inference latency by 21%. Large-scale online A/B tests on Kuaishou’s search platform further reveal notable improvements: CTR +2.15%, CVR +2.27%, OPM +3.60%, Clicks +1.80%, and Orders +3.12%. It demonstrates that OneVision effectively unifies retrieval and personalization, delivering measurable gains while simplifying the traditional serving pipeline.
@@ 2.2. Generative Retrieval System
Generative Retrieval application. Recent advances show that GR models can replace traditional recall–ranking stacks. OneRec unifies recall and ranking via session-wise generation and DPO, achieving stable online gains with simpler serving (Deng et al., 2025). MMQ further learns semantic IDs from multimodal content through a shared–specific tokenizer and behavior-aware fine-tuning to align semantics with user preferences (Xu et al., 2025). In e-commerce search, OneSearch builds an end-to-end system where a keyword-enhanced hierarchical quantization encoder generates semantic IDs combining content, context, and collaborative signals; multi-view user sequences further improve personalization (Chen et al., 2025). These advances show GR can unify fragmented objectives, introduce personalization early, and reduce reliance on hand-crafted indices.
@@ 3.1. Vision-aligned Residual Quantization
A recent approach, RQ-OPQ (Chen et al., 2025), adopts RQ-Kmeans for shallow shared representations, and OPQ for deep unique features, which can encode each item hierarchically and personally. Inspired by this schema, here we propose VRQ, which integrates multi-view contrastive learning for trainable shallow codebooks and leverages statistical business features in residual OPQ encoding to enhance consistency and uniqueness. Additionally, we incorporate category information to suppress noisy samples and improve category-code consistency across the e-commerce image pool.
@@ 3.1.1. Collaborative Image Feature Alignment
In generative vision search systems, aligning similar product images or multi-view images enables the generation of high-quality visual representations, which are crucial for semantic encoding. Firstly, we construct a high-quality dataset of image pairs derived from real e-commerce user search logs, including multi-view images of the same items, query–item, and item–item pairs. Then, we apply cosine similarity for preliminary filtering, employ Qwen-VL (Bai et al., 2023) to further verify semantic consistency, and finally perform manual inspection and refinement to remove noisy data.
@@ 3.1.3. VRQ Hierarchical Quantization Encoding
Concretely, given a single product image and its associated category label, we first extract the visual feature $x$ using the image encoder $\mathcal{V}$ and the category feature $y$ using a pretrained BGE model (Xiao et al., 2024). We then learn latent representations $v:=\mathcal{E}_{v}(x)$ and $t:=\mathcal{E}_{t}(y)$, where $\mathcal{E}_{v}$ and $\mathcal{E}_{t}$ are modality-specific DNN encoders for the visual and textual inputs, respectively. To effectively combine multi-modal features into a unified representation (Baldrati et al., 2023), we compute the fused embedding as
@@
where $f_{i}^{(1)}$ and $f_{i}^{(2)}$ denote fused embeddings of the $i$-th positive item pair, $N$ is the number of positive pairs in a batch, and $\tau$ is a temperature parameter. Using only $\mathcal{L}_{\text{cons}}$ tends to make the model focus solely on visual features. To further enhance the category information (Chen et al., 2023), we introduce a margin loss:
(4) $\mathcal{L}_{\mathrm{mar}}=\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{N}\left[\max\bigl(0,\ -\gamma+s_{ij}^{f2f}-s_{ij}^{v2f}\bigr)\right],$
@@ 3.2.1. Pretraining for Visual–Semantic Alignment
Considering that the basic architectures, such as BART and mT5, are pretrained on large-scale textual corpora, while the input of the vision search system is based on image representations, we first introduce a pretraining stage to establish initial semantic alignment among product images, titles, categories, and SIDs. To this end, we design four related tasks for pretraining: (i) Predicting the product SID from the corresponding image. (ii) Input both image and product title and output the relevant SID. (iii) Predicting the product category from the image. (iv) Input both image and product title and output the corresponding category. The first task focuses on aligning visual representations with SIDs, while the remaining tasks introduce additional modalities to reinforce semantic relevance and facilitate alignment. During unified training, we insert a start token $t_{[BOS]}$ and an ending token $t_{[EOS]}$ at the first and last place, and use a separate token $t_{[SEP]}$ between different modalities. The training loss for the generative model is a cross-entropy loss for next-token prediction, formulated as:
@@ 3.2.3. Personalized Modeling and Post-training
To model user behaviour, we incorporate two complementary, per-user components: (i) user long-term behavioural sequence, represented as a single aggregated SID obtained by fusing the SIDs of multiple products that the user has recently clicked or purchased on the Kuaishou e-commerce platform; and (ii) user short-term behavioural sequence, comprising the individual SIDs of the user’s five most recently clicked items during image-based search. In addition, we introduce explicit search-scene information to better distinguish intent across different contexts. The model is further trained with personalized SFT using the NTP loss to enhance its ability to capture nuanced user-specific purchasing preferences.
@@ 3.3. Dynamic Pruning
While the GR model offers strong semantic understanding, it also introduces significant computational overhead—particularly during beam search decoding. For a single query image, the model must perform $B$ rounds of decoding, where $B$ is the number of beam search candidates. This greatly impacts online inference speed. Moreover, the encoding of image tokens contains redundancy and inefficiency, both in the visual encoder and the generative model’s encoder. To mitigate this, we propose a dynamic pruning strategy that effectively retains only the most informative visual tokens. This significantly reduces inference latency with minimal impact on prediction accuracy for beam search.
@@ 3.4. End-to-End Inference
To produce a well-ranked candidate set, we use beam search. This method maintains multiple SID candidates during generation, selecting them based on the accumulated log-probability scores. Starting with a $t_{[BOS]}$ token, the model predicts the top-$B$ most likely next tokens ($B$ being the beam size), and extends each token sequence. After scoring the extended sequences, only the top-$B$ with the highest cumulative scores are retained. To prevent invalid SIDs, we apply constrained decoding using a Trie structure (De Cao et al., 2020; Fredkin, 1960). The Trie is built from the full set of valid SIDs before inference and restricts decoding to valid prefixes only. This ensures that generated sequences are valid and improves retrieval reliability.
@@ 4.1. Implementation Details
Datasets. We construct training datasets from large-scale industrial e-commerce search logs to support the multi-stage training of OneVision. The corpus consists of three parts: (1) query images and product images, each denoted with its corresponding SIDs, providing discrete representations for both search queries and products; (2) item-to-item (i2i) associations, representing semantically similar product pairs; and (3) 4 months’ data of user search logs from Kuaishou’s platform, covering item click and purchase records. Both the i2i and user-interaction data incorporate multi-view paired samples, enabling the model to build a more comprehensive and robust understanding of each product across diverse visual appearances. For evaluation, we built two test sets. The first is a real-world search interaction set, containing 8.5K click events and 8.5K purchase events sampled from 30 days of historical visual search traffic, designed to measure retrieval quality under actual user behaviours. The second is a semantic benchmark, comprising 67K query–item pairs (about 22K each from top, middle, and tail segments) collected via ANN recall. All datasets are additionally refined using ANN similarity scores and VLM verification to ensure reliable supervision and robust evaluation.
@@ 4.2. Real-World Offline Evaluation
Table 2 presents the ablation study of OneVision. Removing the ranking stage from onlineMCA leads to a noticeable decline in retrieval quality, confirming its importance. Substituting RQ–Kmeans with RQ–OPQ raises HR and MRR improves HR and MRR by 4.04% on average, showing that residual OPQ improves the stability of SIDs and enables a more fine-grained representation. The VRQ with multi-stage training reaches the highest overall scores while remaining comparable to the deployed online system. Adding statistical business characteristics (Convert) information brings an additional 2.98pt in HR and 3.22pt in MRR, which helps the model separate near-duplicate products more effectively. In contrast, user sequence modeling shows limited impact, likely because user preferences are sparse and relatively consistent, offering little additional signal for personalization. Category embeddings mainly mitigate category inconsistency and detection noise, lowering off-category retrievals and strengthening search robustness and user experience, while having a limited direct impact on HR or MRR.